OpenStudy (anonymous):

The position of a kite is given by r(t)=t/8i-3/64t(t-160)j where t>=0 is measured in seconds and distance is measured in meters. At what time does the kite start to lose altitude?

OpenStudy (anonymous):

missing a j?

OpenStudy (anonymous):

theres a j at the end there lol

OpenStudy (anonymous):

The thing is that the i value doesn't influence j (unless there are mountains we don't know about). So we just have y = -3/64t(t-160) which we have to differentiate dy/dt = ?

OpenStudy (anonymous):

that whole part goes together

OpenStudy (anonymous):

We need to know y(t), we don't care about x.

OpenStudy (anonymous):

yea i follow you thanks!

OpenStudy (anonymous):

Can you differentiate?

OpenStudy (anonymous):

lol yes so do i just differentiate the j part with respect to t?

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

thanks then after that i solve for t?

OpenStudy (anonymous):

-3/64(2t - 160) = ?

OpenStudy (anonymous):

i got -3/32(t-80) as the derivative then just solve for t right?

OpenStudy (anonymous):

You need to multiply the single t inside the parentheses At least, that's the easiest way.

OpenStudy (anonymous):

so i got t=5/64 or .078 is that right?

OpenStudy (anonymous):

I'll let you do the arithmetic. The important part is to ignore x, then differentiate correctly and set equal to 0.

OpenStudy (anonymous):

alright i think im good then thank you for your help!